OpenStudy (anonymous):

I am checking my work: The actual measured mass of a beryllium-9 atom is 9.012182 atomic mass units (amu) and the expected mass is calculated to be 9.072429 amu. Therefore, the mass defect for a beryllium-9 nucleus is 0.0624 amu. (1 amu= 1.66X10-27 kg) How much energy will this reaction provide? Answer 3.12 x 10-20 kg m2/s2 9.36 x 10-12 kg m2/s2 3.46 x 10-36 kg m2/s2 1.11 x 10-11 kg m2/s2 I believe it is: 9.36 x 10-12 kg m2/s2?

OpenStudy (unklerhaukus):

∆E = ∆mc^2 ∆m = (9.072429-9.012182) [amu] = 0.060247 [amu] = 0.060247 [amu] * 1.66*10^-27 [kg/amu] = 1.0001002e-28 [kg] c = 2.997925*10^8 [m/s] ∆E = ∆mc^2 = 1.0001002e-28 [kg] * (2.997925*10^8 [m/s])^2 = 8.988452e-12 [kg m^2/s^2] = 8.99e-12 [J]

OpenStudy (unklerhaukus):

close enough

OpenStudy (anonymous):

Thanks so much UnkleRhaukus!!!